OpenStudy (anonymous):

Is (-2, 5) a solution to 2x - 3y = -19 ?

OpenStudy (anonymous):

Am I supposed to plug in those numbers, or?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

that would be the easiest way to check

OpenStudy (khaakenajaf):

subsitute the numbers and the sides will be equal 2(-2)-3(5)=-19 -19=-19